Fermi National Laboratory
 

Accelerator Update for February 23, 2007 - February 26, 2007

 

Friday February 23

The day shift began with Operations monitoring Tevatron (TeV) store 5242 (90.24E30) and the Recycler stash (158.6E10), with the Antiproton Source (Pbar) stacking (37.9E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with the Meson MTest experiment and NuMI taking beam, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:20 AM, a Booster RF station (BRF6) tripped off. It reset okay.

At 8:22 AM, MI-60 suffered a vacuum burst. Many vacuum pumps tripped off. Operations turned them back on.

RF experts changed the bias power supply for MIRF12 at 9:21 AM.

At 10:21 AM, the NuMI beam permit tripped off due to a bad beam position. It reset okay.

At 10:27 AM, MIRF12 tripped off due to a problem with its bias supply. Operators bypassed the station.

At 10:32 AM, Operations reported that part of the MI RF system was down due to a breaker trip. The Duty Electrician was contacted.

Operations began preparing Booster for an access at noon.

Preaccelerator experts accessed the H- Source at 12:30 PM. They were out in 45 minutes.

Operations prepared Linac for an access at 12:41 PM. They were out in an hour.

Linac experts completed their work on LRF5. Linac resumed sending beam at 2:20 PM.

The Booster access ended at 2:34 PM.

At 2:47 PM, BRF7 tripped off for the first of two times in 15 minutes. An expert investigated after the second trip.

Operations resumed sending beam to Pbar, NuMI, and the Meson MTest experiment at 3:08 PM.

Operations reported at 4 PM that the Recycler had made one stash of antiprotons during the day shift.

At 4:52 PM, BRF6 tripped off for the first of five times in 6.5 hours. It reset each time.

The MTest experiment finished using the beam at midnight.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Saturday February 24

The midnight shift began with Operations monitoring store 5242 (35.44E30) and the stash (668.0E10), with Pbar stacking (16.0E10), and with NuMI taking beam.

At 1:27 AM, the NuMI beam permit tripped off due to the cooling fluid temperature for a kicker. This was the first of about eleven trips in less than two hours. It reset each time. Operations notified experts.

At 2:02 AM, BRF6 tripped off. It reset.

At 5:08 AM, the MI beam permit tripped off for the first of three times in 12 minutes due to a thermocouple (I:TC38BB) spiking high. It reset each time.

At 5:24 AM, the MI thermocouple tripped off and would not reset for more than a few seconds before tripping off again. Operations called a MI expert.

Operations reported that at the suggestion of a MI expert they bypassed the thermocouples. There are actually 12 thermocouples associated with this trip and they monitored the temperatures on the MI-8 collimators. The thermocouples were all reading back exactly the same, so the problem was with the readback. Operators masked out the readbacks. Other devices in the tunnel would help monitor what was happening until MI experts could investigate later in the morning.

Operations established beam to MTest at 6:32 AM.

Operations reported at 7:53 AM that the Recycler stashed three times during the midnight shift.

Operations terminated store 5242 at 8:04 AM.

Operations prepared CDF and D0 for access into their collision halls at 8:14 AM.

At 8:29 AM, BRF6 tripped off twice in 25 minutes. It reset.

D0 came out of access at 10:41 AM.

CDF came out of access at 12:21 PM.

Operations put the TeV through a dry squeeze at 12:48 PM.

MTest finished taking beam at 12:53 PM. Their experiment was completed.

Operations put the TeV into shot setup at 1:21 PM.

Operations established store 5244 at 3:41 PM with an initial luminosity of 283.81E30.

At 5:10 PM, NuMI tripped off for the first of seven times in about 2.5 hours on a kicker cooling fluid level alarm. It reset.

At 6:34 PM, MIRF3 tripped off on a filament power supply off indication. Operators investigated and found that the power supply was off. Operators reset it manually. Later it tripped off again and was reset remotely.

Operations reported at 8 PM that the Recycler had stashed four times in the last 12 hours.

At 11:10 PM, Operations reported that they might have had a power glitch. A number of unrelated devices tripped off. Fortunately, the stack, stash, and store weren’t affected. Everything was reset.

Sunday February 25

The midnight shift began with Operations monitoring store 5244 and the stash, with Pbar stacking, and with NuMI taking beam.

At 1:33 AM, Operations reported that there were problems with a Recycler stash and some of the antiprotons were lost during transfers. Operations managed to recover and resume normal stashing.

At 7:13 AM, Operations reported controls problems. Operators could not bring up new MCR pages. Also, the NuMI permit had dropped, but operators couldn’t see why. Operations contacted an expert.

Operations reported at 7:32 AM that they restored normal MCR controls with an ACNET restart command. They resumed sending beam to NuMI.

Operations reported at 8 AM that the Recycler had stashed four times in the last 12 hours.

At 8:38 AM, NuMI tripped off for the first of many times in an hour due to the kicker temperature. It reset each time. A NuMI expert was called in to fix this problem.

At 9:49 AM, Operations halted NuMI beam due to its kicker problem.

Operations resumed sending beam to NuMI at 10:02 AM. The kicker temperature trip level was raised.

Switchyard (SY) experts began a study period at 10:32 AM.

At 11:33 AM, the NuMI horn tripped off twice due to a high temperature indication, but the kicker hadn’t tripped off. The horn reset okay.

At 1:49 PM, the Booster West Anode Power Supply (WAPS) tripped off due to a high temperature indication. Operators investigated and found that the readback had jumped from 87˚F to 117˚F. However, the CUB supply temperature was 85˚F and temperature gauge for BRF12 was at ~84˚F. Operations conferred with a water expert and then bypassed the specific alarm. Other alarms were still active.

SY studies ended at 3:08 PM.

Operations reported at 4:50 PM that LCW experts changed the NuMI kicker water skid filters and the temperatures returned to normal.

Operations terminated store 5244 at 5:30 PM.

Operations put the TeV into shot setup at 6:04 PM.

Operations reported at 8 PM that the Recycler had stashed four times in the last 12 hours.

Operations established store 5245 at 8:50 PM with an initial luminosity of 292.26E30. This store set a New Luminosity Record.

Monday February 26

The midnight shift began with Operations monitoring store 5245 and the stash, with Pbar stacking, and with NuMI taking beam.

At 5:08 AM, the NuMI horn tripped off on a stripline ground fault. It reset.

Operations reported at 8 AM that the Recycler had stashed five times in the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/26/2007    email Fermilab